OpenStudy (anonymous):

The largest swimming pool in the world is in Casablanca. The rectangular pool has a perimeter of 1110 meters, and its length is 30 meters longer than 6 times its width. How many meters wide is it?

OpenStudy (anonymous):

\[P = 1110 \, \textrm{meters} = 2 * (L + W)\] \[L = 6W + 30\] \[W = \frac{L}{6} - 5\] \[1110 \textrm{ meters} = 2 * (6W + 30 + W)\]\[1110 \textrm{ meters} = 14W + 60\]\[W = 75\]

OpenStudy (anonymous):

Isn't it 12w instead of 14w?

OpenStudy (anonymous):

\[6W + 30 + W = 7W + 30\]\[2 * (7W + 30) = 14W + 60\]

OpenStudy (anonymous):

Then, how'd you get 75?

OpenStudy (anonymous):

\[1110 = 14W + 60\]\[1050 = 14W\]\[\frac{1050}{14} = \frac{14W}{14}\]\[75 = W\]

OpenStudy (anonymous):

Sorry, I'm a bit slow. Thank you.

OpenStudy (anonymous):

thank you!!